OpenStudy (anonymous):

The length of a rectangle plus its width is 25cm. The area is 156 square cm. What are the length and width of the rectangle?

OpenStudy (anonymous):

l + w = 25 l * w = 156 Solve the first one for l. Then plug that in for l in the second one and solve for w. Finally, plug that back into the first one to solve for l.

OpenStudy (anonymous):

Well, I had a nice reply, but it didn't post. You could also use a table of values of numbers that add up to 25.

OpenStudy (anonymous):

so l = 25 - w. Sub in: (25 - w)(w) = 156 -w^2 + 25w = 156 -w^2 + 25w - 156 = 0 (w - 12)(-w + 13) = 0 so w is either 12 or 13. if w is 12, l is 13 and vice-versa.

OpenStudy (anonymous):

L W L + W Area 1 24 25 24 2 23 25 46 3 22 25 66 4 21 25 84 5 20 25 100 6 19 25 114 7 18 25 126 8 17 25 136 9 16 25 144 10 15 25 150 11 14 25 154 12 13 25 156 13 12 25 156 14 11 25 154 15 10 25 150 16 9 25 144 17 8 25 136 18 7 25 126 19 6 25 114 20 5 25 100 21 4 25 84 22 3 25 66 23 2 25 46 24 1 25 24